For a gas having molar mass M specific heat at constant pressure canbe given by what in terms of gamma, R and M.??

Grade:12

1 Answers

Krishna Santoshi
5 years ago
Since C= d/dt (3kNA) = 3kNA/ mole = 24.94J/mole.K
since Cp- Cv = R
so, Cp = R + Cv
        Cp = R + 24.94 J/mole.K